OpenStudy (anonymous):

what is the tan(3/4) in degrees?

OpenStudy (anonymous):

use a calculator. i get .01309

OpenStudy (anonymous):

it is cot(4/3)

OpenStudy (callisto):

my calculator tells me it is 0.01309 ....

OpenStudy (anonymous):

I assume you mean arctan(3/4) as tan(3/4) would not result in an angle. arctan(3/4)=36.87

OpenStudy (anonymous):

thanks everyone

OpenStudy (anonymous):

hold the phone. are you looking for the angle whose tangent is 3/4 ?

OpenStudy (anonymous):

no i was looking to see what the tan(3/4) would be in degrees bt now i use the arctan in order to get the answer of tan(3/4) in the degrees.

OpenStudy (anonymous):

tan(pi/4) = 1 radian where pi is approx 3.14159, so the tan(3/4) is going to be just under 1 rad. To convert radians to degrees you divide 180/pi. Thus tan(3/4) is approx 0.932 and 0.932*(180/pi) = 53.38 degrees.

OpenStudy (anonymous):

\[\text{Tan}[3/4]=\text{Tan}\left[\frac{3}{4}\frac{180}{\pi }\right]=\text{Tan}\left[\frac{135}{\pi }{}^{\circ}\right]=\text{Tan}[42.9718{}^{\circ}] \]